OpenStudy (anonymous):

9.01], [9.02] Part 1: Solve each of the quadratic equations below. Show your work. (3 points) x2 – 36 = 0 and x2 = 8x – 12 Part 2: Describe what the solution(s) represent to the graph of each. (2 points) Part 3: How are the graphs alike? How are they different? (2 points)

OpenStudy (anonymous):

I really don't know this one. :)

OpenStudy (anonymous):

\[x ^{2}-36\rightarrow x ^{2}=36\rightarrow \sqrt{x ^{2}}=\pm \sqrt{{36}}\]

OpenStudy (anonymous):

Thx, the other one?

OpenStudy (anonymous):

\[x=\pm6\]

OpenStudy (anonymous):

still the first one the other will be....

OpenStudy (anonymous):

\[x ^{2}=8x-12\rightarrow x ^{2}-8x+12=0\rightarrow (x-2)(x-6)=0\]\[x=2, x=6\]

OpenStudy (anonymous):

solutions are where the graph crosses the x-axis

OpenStudy (anonymous):

The graphs are alike in that they are both parabolas and they each cross the x-axis at x=6 they differ in that the other x-intercepts are not the same.

OpenStudy (anonymous):

Thanks so much! :D

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):